OpenStudy (anonymous):

40.32 is 76.8% of what number?

OpenStudy (gorv):

let it is x%

OpenStudy (gorv):

\[127.5=51*\frac{ x }{ 100 }\]

OpenStudy (gorv):

multiply by 100 on both side

OpenStudy (anonymous):

12700 & 5100?

OpenStudy (gorv):

127.5*100=51*x

OpenStudy (gorv):

on right side 100 will get cancelled by 100 in denominator

OpenStudy (anonymous):

so it will be 12700=51x?

OpenStudy (gorv):

127.5*100=12750

OpenStudy (anonymous):

now How do we know the percent?

OpenStudy (gorv):

12750=51x divide both side by 51

OpenStudy (gorv):

\[\frac{ 12750 }{ 51 }=x\]

OpenStudy (anonymous):

250

OpenStudy (anonymous):

what about the B question

OpenStudy (gorv):

yeah 250 %

OpenStudy (gorv):

let the number is x

OpenStudy (gorv):

40.32=x*76.8/100

OpenStudy (anonymous):

then i multyply by 100 on the both side ?

OpenStudy (gorv):

yeah :)

OpenStudy (anonymous):

ok thank you